OpenStudy (anonymous):

7/3 minus 3/2 plus 3/8 how do I come up with a common denominator?

OpenStudy (anonymous):

Multiply your 3 and your 8. You get 24. So that will work for your 7/3, your 3/8, and of course your 3/2 as well.

OpenStudy (anonymous):

if you multiply 7/3 by 8/8, 3/8 by 3/3, and 3/2 by 12/12, see what happens

OpenStudy (anonymous):

You can do that because each time you're only multiplying by 1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

\[\frac{7}{3}- \frac{3}{2} + \frac{3}{8}=\frac{29}{24} \]